Graph the equation and identify the specified parts (x-2)^2/9+(y-1)^2/25=1?
Find the
Vertices:
Co- vertices:

Answer by lwsshak3(11628)

Graph the equation and identify the specified parts (x-2)^2/9+(y-1)^2/25=1?
Find the
Vertices:
Co- vertices:
..
Standard form of ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1, (a>b), with (h,k) being the (x,y) coordinates of the center.
Standard form of ellipse with vertical major axis: (x-h)^2/b^2+(y-k)^2/a^2=1, (a>b), with (h,k) being the (x,y) coordinates of the center.
The difference between the two forms is the interchange of a^2 and b^2.
..
(x-2)^2/9+(y-1)^2/25=1
Center at (2,1), with a vertical major axis.(Second form listed above)
..
a^2=25
a=5
length of major axis=2a=10
end points of major axis, (2,1±5)
..
b^2=9
b=3
length of co-vertices or minor axis=2b=6
end points of minor axis, (2±3,1)
see graph below as a visual check on answers:
note the center and lengths of the minor and major axes and their end points
..
y=(25-25(x-2)^2/9)^.5+1